OpenStudy (anonymous):

How much H3PO4 is needed to neutralize 1 liter of a 0.75M solution of NaOH?

OpenStudy (cuanchi):

1/3 liter of 0.75M H3PO4

OpenStudy (cuanchi):

1 liter of 0.25 M H3PO4

OpenStudy (rushwr):

3NaOH + H3PO4 ---> Na3PO4 + 3H2O As we can see 3 moles of NaOH reacts with 1 mole of H3PO4 . 1st find the no. of moles of NaOH use n =cv n= no. of moles C = concentration V = Volume n= 1* 0.75 = 0.75 moles 3 moles of NaOH will react with 1 mole of H3PO4 so how much of H3PO4 will react with 0.75 moles of NaOH. that would be 075/3 = 0.25 moles of H3PO4